OpenStudy (anonymous):

2x^2=5x solve the equation

OpenStudy (anonymous):

2x^2 - 5x = 0 x(2x -5) = 0 x = 0 2x-5 = 0 => x = 5/2

OpenStudy (anonymous):

Okay This can be solve using completing the square technique. 2x^2=5x 2x^2-5x =0 2x^2-5x +( Some number )=0+( Some number ) Right Now we can find out that (Some number ) by using the (-b/2a)^2 here, our b is -5, a is 2 ( from the form ax^2+bx+c =0) (-b/2a)^2 = (5/4)^2 2x^2 -5x +( (5/4)^2)= (5/4)^2 Hope you can soleve the LHS now If you get (x+A)(x+B) =(5/4)^2, u know, (x+A) =(5/4)^2 or (x+B)=(5/4)^2 Now u know how to solve for X. Aske me if it is not clear.